# Fleurs-SLU: A Massively Multilingual Benchmark for Spoken Language Understanding

### v2

@@ 4.3 Training Setup
Training Details. We train all models with AdamW (Loshchilov and Hutter, 2019), weight decay of , and with 10% linear warm-up and followed by linear decay on an effective batch size of 32. For SIB-Fleurs, we run a grid search over the learning rates , since suitable hyperparameters have not yet been extensively studied for mSE on such a downstream task. For Belebele-Fleurs, we fine-tune the LLMs on the learning rates with LoRAs of rank and alpha attached onto all linear layers, as full fine-tuning is prohibitively expensive. We train models for 20 and 3 epochs for SIB-Fleurs and Belebele-Fleurs, respectively, and validate at every 10% of training steps. Experimental results are averaged across 3 random seeds. We report results for runs on the learning rate that performs best, on average, on the English validation sets.
@@
Text vs. Utterance Quality. We evaluate models trained on the original text (Text) as well as on varying quality of speech. For each sentence, Fleurs comprises one or more utterances. We compute the character error rate (CER) between the ground truth sentence and its transcriptions.111111We compute CER against the original Flores sentence, as Flores represents the written form of the utterance best suited for processing by LLMs. We then fine-tune models on two sets of utterances: those with the lowest CER (best quality) and those with the highest CER (worst quality). We thereby isolate the downstream impact of utterance quality, ensure a balanced training set, and compare mSE to CS more fairly, which benefit less from mixing transcriptions of varying quality.
@@ 5 Results
In-Language ZS-XLT Utterance Quality Best Worst Best Worst Best Worst Best Worst Best Worst SIB-Fleurs EN (1) Whisper (85) SeamlessM4T (90) Unsupported (7) Not-EN (101) mSE mHuBERT 41.1 39.7 27.3 19.1 27.2 19.1 26.5 16.3 26.9 18.6 MMS-1b 44.6 46.9 19.9 13.6 19.8 13.5 20.8 13.3 19.8 13.4 MMS-1b-Fleurs 64.8 64.2 25.2 20.8 24.7 20.3 27.1 22.1 25.0 20.6 MMS-1B-All 55.2 53.1 21.7 18.1 21.4 17.9 24.9 21.0 21.7 18.2 Whisper-v3-L 78.3 76.3 46.0 42.0 45.5 41.3 39.4 38.4 44.7 40.8 SeamlessM4Tv2-L 87.4 88.5 82.3 79.3 82.2 79.1 55.7 52.9 79.0 75.8 CS Roberta-L-Wh-EN 90.4 91.7 75.8 72.7 73.6 70.8 52.0 48.7 71.5 68.4 Roberta-L-S4T-EN 91.7 90.8 86.2 84.5 85.9 84.2 56.3 55.2 82.3 80.6 LLM2Vec-L-Wh-EN 91.1 92.3 76.6 75.3 74.5 73.4 53.0 53.2 72.4 71.1 LLM2Vec-L-S4T-EN 93.0 91.3 87.1 85.2 86.9 84.9 58.1 56.6 83.4 81.4 NLLB-LLM2Vec-L-Wh 92.5 91.0 77.5 73.9 75.4 72.2 54.4 53.2 73.7 70.4 NLLB-LLM2Vec-L-S4T 94.0 92.7 85.1 84.9 85.3 84.9 62.3 60.7 82.3 81.9 Text Roberta-L-S4T-EN 92.3 87.4 89.1 55.3 85.2 LLM2Vec-S4T-EN 91.0 86.4 87.9 55.6 84.1 NLLB-LLM2Vec 92.3 88.1 87.9 80.7 87.2 Belebele-Fleurs EN (1) Whisper (79) SeamlessM4T (84) Unsupported (4) Not-EN (91) CS Roberta-L-Wh-EN 82.6 81.3 57.5 56.2 56.4 55.2 46.2 44.7 55.5 54.3 Roberta-L-S4T-EN 82.0 79.4 67.5 66.4 67.0 65.8 47.2 46.2 65.2 64.1 LLM2Vec-Wh-EN 94.9 94.4 74.6 73.5 73.6 72.6 63.5 62.9 72.7 71.8 LLM2Vec-S4T-EN 95.5 94.4 84.0 83.1 83.5 82.5 64.7 63.6 81.8 80.9 NLLB-LLM2Vec-Wh 94.8 93.5 58.0 56.2 56.8 55.0 40.9 41.0 55.5 53.8 NLLB-LLM2Vec-S4T 94.9 93.5 61.3 60.2 60.8 59.6 43.9 41.1 59.2 57.8 Text Roberta-L-S4T-EN 83.3 72.1 72.6 48.3 70.5 LLM2Vec-S4T-EN 95.3 87.8 88.3 66.0 86.3 NLLB-LLM2Vec 95.1 65.6 64.6 55.2 63.8
@@
English. The first column presents the in-language English performance by task, modality (Text vs. mSE and CS), and utterance quality by model.
@@
Text & CS. English ASR performance is strong for both transcription models (cf. Figure 1), and CS effectively leverages the NLU capabilities of LLMs, thanks to their extensive pre-training on both tasks. CS are even on par with models trained on gold text, underscoring the quality of ASR on English. As a result, CS outperforms all mSE models on SIB-Fleurs. On Belebele-Fleurs, both CS models backed by LLM2Vec and NLLB-LLM2Vec perform comparably. In sum, the transcription model has no clear impact on English performance.
@@
mSE. The English results on SIB-Fleurs suggest that the SLU capabilities of mSE models are strongly shaped by their pre-training curriculum. mHubert (41.1 %) and MMS-1b (44.6%), which are pre-trained solely in a self-supervised manner, underperform other mSE on SIB-Fleurs. Fine-tuning MMS-1b on ASR, either on Fleurs (MMS-1B-Fleurs) or Fleurs combined with additional data (MMS-1b-All), results in significant improvements on SIB-Fleurs of +10-20 percentage points (pp). Notably, MMS-1B-Fleurs outperforms MMS-1b-All (+9.6 pp), suggesting that the broader domain mixture in the MMS-1b-All training set negatively affects this task. Whisper-v3-Large surpasses all MMS-1B variants (78.3%). The large-scale pre-training of Whisper-v3 on both multilingual ASR and S2ETT enhances its SLU performance. SeamlessM4Tv2-Large is the best performing mSE on the English test set of SIB-Fleurs (87.4%), with only a slight performance gap compared to CS (ca. -3 pp). The joint pre-training on multilingual ASR as well as multilingual and cross-modal MT with text-to-speech knowledge distillation (KD) significantly enriches the semantics in speech representations of SeamlessM4Tv2-Large. The results particularly highlight that (i) multilingual ASR, (ii) S2ETT pre-training, and (iii) text-to-speech KD are crucial training objectives for enabling speech encoders to acquire strong SLU capabilities.
@@
Utterance Quality. The quality of English utterances does not affect performance for either mSE or CS across both tasks. This can be attributed to the exhaustive training of all models on English ASR due to the scale of English speech training data. This is why CS are also on par with Text.
@@
zs-xlt. The right-hand side of Table 3 reports zs-xlt performance. For each task, we group the languages into (i) languages supported by Whisper-v3, (ii) languages supported by SeamlessM4Tv2, and (iii) languages unsupported by either model. Whisper-v3 and SeamlessM4Tv2 overlap in their support for 81 Fleurs languages.121212For unsupported languages, we hand-select the closest supported language to transcribe into.
@@
Text & CS. Across both tasks and all language groups, models trained on transcriptions of SeamlessM4Tv2-Large consistently outperform models fine-tuned on transcriptions of Whisper-v3-Large. The performance gap grows as Whisper’s support decreases or becomes unavailable for the target languages (cf. ‘SeamlessM4Tv2’ and ‘unsupported’ language groups in Table 3). Nevertheless, while CS are mostly competitive on SIB-Fleurs, they trail models evaluated on ground-truth paragraphs more significantly on Belebele-Fleurs (ca. -5 pp). We presume that transcription and S2ETT errors propagate more severely in multiple-choice QA. For SIB-Fleurs, all CS with SeamlessM4Tv2-Large deteriorate only slightly in xlt performance to all 101 target language, on average, relative to English (ca. -10 pp). Most of this xlt gap comes from the 7 target languages that SeamlessM4Tv2 does not support (ca. -35 pp on average). In contrast, pairing LMs with Whisper-v3-Large causes more pronounced drops on languages that Whisper supports (ca. -13 pp) and again a larger deficit on languages neither model supports (ca. -32 pp). CS are highly competitive to Text (ca. -0.7 pp). Only NLLB-LLM2Vec evaluated on gold text tremendously outperforms all other models on unsupported languages, since they are supported by NLLB Team et al. (2022). For Belebele-Fleurs, S2ETT paired with LLM2Vec (i.e., TTest) outperforms zs-xlt cascading on in-language transcriptions and NLLB-LLM2Vec, except for unsupported languages. This suggests that S2ETT of both Whisper-v3 and SeamlessM4Tv2 sufficiently translates the target languages into English for successful NLU. On unsupported languages, however, NLLB-LLM2Vec performs better for likely two reasons. First, for languages with low S2ETT quality, in-language transcriptions to closely related languages likely better preserve the core meanings of input sequences. Second, in addition to translation, NLLB was pre-trained with denoising autoencoding, making NLLB-LLM2Vec more resistant to noisy inputs than LLM2Vec. Overall, the findings suggest that SeamlessM4Tv2-Large is more robust for both in-language ASR as well as S2ETT (cf. Figure 1).
@@
MSE. The zs-xlt performance of mSE on SIB-Fleurs mirrors the trends we observed for English performance. When trained solely with the wav2vec-BERT 2.0 objective, MMS-1B (19%) only slightly surpasses random performance (14%) in zs-xlt for the highest-quality utterances, regardless of the target language group. In contrast to English, post-hoc modular ASR fine-tuning of MMS-1B with language adapters and language-specific decoding heads, whether on Fleurs alone (MMS-1B-Fleurs) or on Fleurs with additional data (MMS-1B-All), yields much smaller gains in zs-xlt (ca. +1-3 pp). Furthermore, despite employing fully parameter-shared multilingual ASR and S2ETT pre-training, Whisper-v3-Large fails to transfer strong English in-language performance to other languages effectively, with the xlt gap ranging from -32.3 pp for supported languages to -38.9 pp for unseen languages. In contrast, only SeamlessM4Tv2-Large achieves performance comparable to English on supported languages (-5.2 pp) though also deflates in performance on unsupported languages (55.7%). Overall, these findings nevertheless indicate that multilingual cross-modal translation and multilingual text-to-speech distillation align and semantically enrich multilingual speech representations to enhance cross-lingual SLU.
@@
Utterance Quality. The utterance quality consistently affects performance on both tasks for all model configurations in zs-xlt. Notably, CS seem to be less affected than mSE. We attribute this observation to the much more sizable pre-training on text of diverse quality of LLMs compared to speech models. For SIB-Fleurs, the magnitude of the performance gaps between best and worst quality utterances in zs-xlt mimics how well the speech model backbones perform across both tasks in both CS and end-to-end speech classification. SeamlessM4Tv2-Large is more robust to noisy utterances than Whisper-v3-Large, whereas MMS variants suffer from the largest drops (ca. 5 pp).
@@ 5.1 Further Analyses and Discussion
ASR & Translation Performance. To further understand the underlying factors behind our main results (cf. Table 3), we benchmark Whisper-v3-Large and SeamlessM4Tv2-Large on ASR and Speech-to-English-text-translation (S2ETT) on all 102 Fleurs languages. We first compute CER and sacreBLEU between the Flores sentences and the in-language transcriptions and S2ETT transcriptions of utterances in Fleurs pooled across all splits, respectively.131313We verify that there is no leakage of dev and test splits of Fleurs in SeamlessM4Tv2-Large and Whisper-v3-Large. We observe that the averages and the standard deviations CER and sacreBLEU by task and language for either model is highly comparable across splits. We trust that the models have not been trained on the dev and test set of Fleurs. We use Huggingface’ evaluate to compute for CER and sacreBELU. We then ‘macro-average’ the metrics over all languages. Figure 1 charts the results of this analysis.
@@
S2ETT. SeamlessM4Tv2-Large outperforms Whisper-v3-large across the board in speech-to-English-text-translation and is favored for all but 2 languages. For both groups of languages supported by, Whisper-v3-Large and SeamlessM4Tv2-Large, respectively, SeamlessM4Tv2-Large achieves on average about 8 higher sacreBLEU than Whisper-v3-Large. The gap from SeamlessM4T-Large to Whisper-v3-Large reduces to about 5 sacreBLEU for unsupported languages. This supports the notion that SeamlessM4Tv2-Large is a more robust multilingual speech encoder.
The more robust ASR and much stronger translation performance of SeamlessM4Tv2-Large results stem from its pre-training. SeamlessM4Tv2 first initializes a text encoder and decoder with weights from a pre-trained NLLB translation model Team et al. (2022) and a speech encoder pre-trained on 4.5M hours of self-supervised training on wav2vec-BERT 2.0 objective. The model is then trained on translation objectives from text and speech to text between any two languages in both translation directions. The text encoder-decoder backbone is used to train the speech encoder with token-level knowledge distillation objectives on decoder output representations. On the contrary, Whisper-v3 trains models from scratch on, among others, in-language ASR and S2ETT (cf. §4). Consequently, the mixture of strong initialization from existing MT backbones, text-to-speech knowledge distillation, and multimodal translation objectives result in much stronger translation performance.
@@
Length Adaptor. The sequence length of encoded speech typically far exceeds the length of embedded tokenized text for the same input. The mSE of SeamlessM4Tv2-Large thus appends a temporal convolution, a ‘length adaptor’, as its final layer to reduce the resolution of speech frames by a factor of 8 and to better align the modalities Seamless Communication et al. (2023).141414For architectural details on the length adaptor, we refer to the original paper Communication et al. (2023).
This begs the question whether a length adaptor is essential for pooling semantics from speech output tokens. We hence compare SeamlessM4Tv2-Large’s performance on SIB-Fleurs with and without the length adaptor.
Table 4 presents the inconclusive results for both variants. For English, performance improves on high-quality utterances but decreases on lower-quality ones. In contrast, performance most pronouncedly declines for low-quality utterances in zs-xlt to supported languages (-2.6%). The zs-xlt unsupported languages is not affected for high-quality utterances (-0.1%) and only slightly for low-quality ones (-0.9%). Three factors may explain this modest gap. First, removing the length adaptor reduces model size by 47M parameters (-7.4 pp), which were part of substantial pre-training. Second, the length adaptor, as the final layer of the mSE, is most explicitly trained to embed speech into a shared multilingual space, attended to by the text decoder at every layer. Finally, the length adaptor might filter noisy frames through temporal downsampling to improve the robustness of speech encodings. While a modest gap persists across setups, we cannot decisively infer whether appending a length adaptor is, as opposed to replacing model capacity with other layers of equal parameter size during pre-training, crucial for improved multilingual SLU. We conclude that the pre-training regime is more vital for multilingual SLU capabilities of mSE than nuanced architectural design choices.
ZS-XLT EN (1) S4T (89) N/A (12) AVG (101) S4Tv2-L Size B W B W B W B W Incl. LA 635M 87.4 88.5 82.2 79.1 55.7 52.9 79.0 75.8 Excl. LA 588M 88.5 86.4 80.9 77.5 55.6 52.0 77.8 74.4 47M +1.1 -2.1 -1.3 -2.6 -0.1 -0.9 -1.2 -1.4

### v3

@@ 4.3 Further Details
Hyperparameters. We train all mSE and CS models with AdamW (Loshchilov & Hutter, 2019), weight decay of , and with 10% linear warm-up and followed by linear decay on an effective batch size of 32. For SIB-Fleurs, we run a grid search over the learning rates , since suitable hyperparameters have not yet been extensively studied for mSE on such a downstream task. For Belebele-Fleurs, we fine-tune the LLMs on the learning rates with LoRAs of rank and alpha attached onto all linear layers, as full fine-tuning is prohibitively expensive. We train models for 20 and 3 epochs for SIB-Fleurs and Belebele-Fleurs, respectively, and validate at every 10% of training steps. Experimental results are averaged across 3 random seeds. We report results for runs on the learning rate that performs best, on average, on the English validation sets. For experiments with Qwen 2.5 7B-Omni and Gemini 2-Flash, we perform greedy decoding. We report further experimental details of our prompting experiments in §A.1.
@@
Text vs. Utterance Quality. We evaluate models trained on the original text (Text) as well as on speech data of varying quality. Each sentence in Fleurs is associated with one or more utterances. To quantify speech quality, we compute the character error rate (CER) between the reference Flores sentence and its transcriptions produced by Whisper-v3-Large and SeamlessM4Tv2-Large. Based on these CER scores, we construct two utterance subsets: one comprising the lowest-CER utterances (best quality), and one with the highest-CER utterances (worst quality). We fine-tune the mSE and CS models and evaluate all models (incl. speech-LLMs) separately on each subset to isolate the effect of utterance quality on downstream performance. This setup also balances the training data to enable a fair comparison between mSE, CS, and Text.
@@ 5 Results
English. The first column presents the in-language English performance by task, modality (Text vs. mSE, CS, and speech-LLM), and utterance quality by model.
@@
Text, CS, and Speech-LLMs. English ASR performance is strong for both transcription models (cf. Figure 1), and CS effectively leverages the NLU capabilities of LLMs, thanks to their extensive pre-training on both tasks. CS are even on par with and sometimes outperform models trained on gold text, underscoring the quality of ASR on English. The slight outperformance of CS over Text in SIB-Fleurs results from model selection on comparatively small validation splits (cf. Table 2). As a result, CS outperforms all mSE models on SIB-Fleurs. On Belebele-Fleurs, both CS models backed by LLM2Vec and NLLB-LLM2Vec perform comparably. In sum, the transcription model has no clear impact on English performance. The speech-LLMs perform on par or better on Belebele-Fleurs while slightly trailing the best fine-tuned models on both CS and Text on SIB-Fleurs. Gemini 2.0 Flash consistently performs better than Qwen 2.5B 7B-Omni. Notably, their cross-modal gap (Text vs. speech-LLM) is small (approx. -1.5%).
English Non-English Utterance Quality (Best; Worst) B W B W B W B W B W Language Groups (Size) Setup EN Whisper S4T Unsup. Non-EN SIB-Fleurs (1) (85) (90) (7) (101) mSE mHuBERT X 41.1 39.7 27.3 19.1 27.2 19.1 26.5 16.3 26.9 18.6 MMS-1b X 44.6 46.9 19.9 13.6 19.8 13.5 20.8 13.3 19.8 13.4 MMS-1b-Fleurs X 64.8 64.2 25.2 20.8 24.7 20.3 27.1 22.1 25.0 20.6 MMS-1b-All X 55.2 53.1 21.7 18.1 21.4 17.9 24.9 21.0 21.7 18.2 Whisper-v3-L X 78.3 76.3 46.0 42.0 45.5 41.3 39.4 38.4 44.7 40.8 SeamlessM4Tv2-L X 87.4 88.5 82.3 79.3 82.2 79.1 55.7 52.9 79.0 75.8 CS Roberta-Wh-EN T 90.4 91.7 75.8 72.7 73.6 70.8 52.0 48.7 71.5 68.4 Roberta-S4T-EN T 91.7 90.8 86.2 84.5 85.9 84.2 56.3 55.2 82.3 80.6 LLM2Vec-Wh-EN T 91.1 92.3 76.6 75.3 74.5 73.4 53.0 53.2 72.4 71.1 LLM2Vec-S4T-EN T 93.0 91.3 87.1 85.2 86.9 84.9 58.1 56.6 83.4 81.4 NLLB-LLM2Vec-Wh X 92.5 91.0 77.5 73.9 75.4 72.2 54.4 53.2 73.7 70.4 NLLB-LLM2Vec-S4T P 94.0 92.7 85.1 84.9 85.3 84.9 62.3 60.7 82.3 81.9 Speech- Qwen 2.5 7B-Omni P 87.0 87.6 59.5 58.4 58.3 57.3 45.6 45.6 56.9 55.9 LLM Gemini 2.0 Flash P 87.6 88.7 82.1 81.4 80.8 80.0 59.4 59.3 78.6 77.8 Text Roberta-S4T-EN T 92.3 87.4 89.1 55.3 85.2 LLM2Vec-S4T-EN T 91.0 86.4 87.9 55.6 84.1 NLLB-LLM2Vec X 92.3 88.1 87.9 80.7 87.2 Qwen 2.5 7B-Omni P 88.7 76.4 75.1 53.9 73.0 Gemini 2.0 Flash P 90.4 87.0 86.8 77.2 86.1 Belebele-Fleurs (1) (79) (84) (4) (91) CS Roberta-Wh-EN T 82.6 81.3 57.5 56.2 56.4 55.2 46.2 44.7 55.5 54.3 Roberta-S4T-EN T 82.0 79.4 67.5 66.4 67.0 65.8 47.2 46.2 65.2 64.1 LLM2Vec-Wh-EN T 94.9 94.4 74.6 73.5 73.6 72.6 63.5 62.9 72.7 71.8 LLM2Vec-S4T-EN T 95.5 94.4 84.0 83.1 83.5 82.5 64.7 63.6 81.8 80.9 NLLB-LLM2Vec-Wh X 94.8 93.5 58.0 56.2 56.8 55.0 40.9 41.0 55.5 53.8 NLLB-LLM2Vec-S4T X 94.9 93.5 61.3 60.2 60.8 59.6 43.9 41.1 59.2 57.8 Speech- Qwen 2.5 7B-Omni P 91.6 91.0 51.4 50.8 50.4 49.7 34.4 34.4 49.0 48.4 LLM Gemini 2.0 Flash P 94.1 93.4 84.4 83.5 82.9 82.0 59.7 59.4 81.1 80.2 Text Roberta-S4T-EN T 83.3 72.1 72.6 48.3 70.5 LLM2Vec-S4T-EN T 95.3 87.8 88.3 66.0 86.3 NLLB-LLM2Vec X 95.1 65.6 64.6 55.2 63.8 Qwen 2.5 7B-Omni P 94.1 69.1 67.4 36.6 64.9 Gemini 2.0 Flash P 95.9 90.8 89.7 82.0 89.1
@@
mSE. The English results on SIB-Fleurs suggest that the SLU capabilities of mSE models are strongly shaped by their pre-training curriculum. mHubert (41.1%) and MMS-1B (44.6%), which are pre-trained solely in a self-supervised manner, underperform other mSE on SIB-Fleurs. Fine-tuning MMS-1B on ASR, either on Fleurs (MMS-1B-Fleurs) or Fleurs combined with additional data (MMS-1B-All), results in significant improvements on SIB-Fleurs of +10-20 percentage points (pp). Notably, MMS-1B-Fleurs outperforms MMS-1B-All (+9.6 pp), suggesting that the broader domain mixture in the MMS-1B-All training set negatively affects this task. Whisper-v3-Large surpasses all MMS-1B variants (78.3%). The large-scale pre-training of Whisper-v3 on both multilingual ASR and S2ETT enhances its SLU performance. SeamlessM4Tv2-Large is the best performing mSE on the English test set of SIB-Fleurs (87.4%), with only a slight performance gap compared to CS (approx. -3 pp). The joint pre-training on multilingual ASR as well as multilingual and cross-modal MT with text-to-speech knowledge distillation (KD) significantly enriches the semantics in speech representations of SeamlessM4Tv2-Large. The results particularly highlight that (i) multilingual ASR, (ii) S2ETT pre-training, and (iii) text-to-speech KD are crucial training objectives for enabling speech encoders to acquire strong SLU capabilities.
@@
Utterance Quality. The quality of English utterances does not affect performance of the models across both tasks. This can be attributed to the large-scale training on English ASR data of all models. This is why CS are also on par with Text.
@@
zs-xlt. The right-hand side of Table 3 reports zs-xlt performance. For each task, we group the languages into (i) languages supported by Whisper-v3, (ii) languages supported by SeamlessM4Tv2, and (iii) languages unsupported by either model. Whisper-v3 and SeamlessM4Tv2 overlap in their support for 81 Fleurs languages.131313For unsupported languages, we hand-select the closest supported language to transcribe into.
@@
Text, CS, and Speech-LLMs. Across both tasks and all language groups, models trained on transcriptions of SeamlessM4Tv2-Large consistently outperform models fine-tuned on transcriptions of Whisper-v3-Large. The performance gap grows as Whisper’s support decreases or becomes unavailable for the target languages (cf. ‘S4T’ and ‘Unsup.’ language groups in Table 3). Nevertheless, while CS are mostly competitive on SIB-Fleurs, they trail models evaluated on ground-truth paragraphs more significantly on Belebele-Fleurs (approx. -5 pp). We presume that transcription and S2ETT errors propagate more severely in multiple-choice QA. Moreover, the best speech-LLM, Gemini 2.0 Flash, performs highly competitively across tasks, boasting on par or better performance than other models.
For SIB-Fleurs, all CS with SeamlessM4Tv2-Large deteriorate only slightly in xlt performance to all 101 target language, on average, relative to English (approx. -10 pp). Most of this xlt gap comes from the 7 target languages that SeamlessM4Tv2 does not support (approx. -35 pp on average). In contrast, pairing LMs with Whisper-v3-Large causes more pronounced drops on languages that Whisper supports (approx. -13 pp) and again a larger deficit on languages neither model supports (approx. -32 pp). CS are highly competitive to Text (approx. -0.7 pp). Only NLLB-LLM2Vec evaluated on gold text tremendously outperforms all other models on unsupported languages, since they are supported by NLLB (Team et al., 2022).
For Belebele-Fleurs, S2ETT paired with LLM2Vec (i.e., TTest) outperforms zs-xlt cascading on in-language transcriptions and NLLB-LLM2Vec, except for unsupported languages. This suggests that S2ETT of both Whisper-v3 and SeamlessM4Tv2 sufficiently translates the target languages into English for successful NLU. On unsupported languages, however, NLLB-LLM2Vec performs better for likely two reasons. First, for languages with low S2ETT quality, in-language transcriptions to closely related languages likely better preserve the core meanings of input sequences. Second, in addition to translation, NLLB was pre-trained with denoising autoencoding, making NLLB-LLM2Vec more resistant to noisy inputs than LLM2Vec. Overall, the findings suggest that SeamlessM4Tv2-Large is more robust for both in-language ASR as well as S2ETT (cf. Figure 1).
The performance of speech-LLMs varies substantially across models and language groups. Gemini 2.0 Flash strongly outperforms Qwen 2.5 7B-Omni on both tasks in non-English settings. Qwen exhibits a marked performance degradation when transitioning from English to non-English inputs, dropping by approx. 30 pp in Text, and even more steeply in the speech modality, where the cross-modal gap widens from approx. 3% (English) to 18% (non-English). In contrast, Gemini maintains strong performance on non-English Text inputs and shows only a moderate cross-modal drop of approx. 9%. This however exceeds the gap for the best fine-tuned CS models vs. their Text counterparts (approx. 3%).
Qwen relies on Whisper-v3-Large as its audio encoder, making it sensitive to Whisper’s transcription limitations. Moreover, its instruction tuning possibly is focused on a small subset of high-resource languages, potentially limiting its generalization to languages not seen during alignment or fine-tuning (QwenOmni). This is consistent with its sharp drop in performance on unsupported languages. By contrast, Gemini’s closed-source nature precludes detailed analysis, but its robustness across modalities and languages suggests a more extensive and diverse instruction tuning setup. It may also benefit from large-scale multilingual pretraining, enhanced alignment objectives, or broader training coverage.
@@
mSE. The zs-xlt performance of mSE on SIB-Fleurs mirrors the trends we observed for English performance. When trained solely with the wav2vec 2.0 objective, MMS-1B (19%) only slightly surpasses random performance (14%) in zs-xlt for the highest-quality utterances, regardless of the target language group. In contrast to English, post-hoc modular ASR fine-tuning of MMS-1B with language adapters and language-specific decoding heads, whether on Fleurs alone (MMS-1B-Fleurs) or on Fleurs with additional data (MMS-1B-All), yields much smaller gains in zs-xlt (approx. +1-3 pp). Furthermore, despite employing fully parameter-shared multilingual ASR and S2ETT pre-training, Whisper-v3-Large fails to transfer strong English in-language performance to other languages effectively, with the xlt gap ranging from -32.3 pp for supported languages to -38.9 pp for unseen languages. In contrast, only SeamlessM4Tv2-Large achieves performance comparable to English on supported languages (-5.2 pp), though also deflates in performance on unsupported languages (55.7%). Overall, these findings nevertheless indicate that multilingual cross-modal translation and multilingual text-to-speech distillation align and semantically enrich multilingual speech representations to enhance cross-lingual SLU.
@@
Utterance Quality. The utterance quality consistently affects performance on both tasks for all model configurations evaluated on non-English languages. Notably, CS and speech-LLMs seem to be less affected than mSE. We attribute this to the much more sizable pre-training on text of diverse quality of LLMs compared to speech-only models. For SIB-Fleurs, the magnitude of the performance gaps between best and worst quality utterances in zs-xlt mimics how well the speech model backbones perform across both tasks in both CS and end-to-end speech classification. SeamlessM4Tv2-Large is more robust to noisy utterances than Whisper-v3-Large, whereas MMS variants suffer from the largest drops (approx. 5 pp).
@@ 5.1 Further Analyses and Discussion
ASR & Translation Performance. To further understand the underlying factors behind our main results (cf. Table 3), we benchmark Whisper-v3-Large and SeamlessM4Tv2-Large on ASR and Speech-to-English-text-translation (S2ETT) on all 102 Fleurs languages in Figure 1. We first compute CER and sacreBLEU between the Flores sentences and the in-language transcriptions and S2ETT transcriptions of utterances in Fleurs pooled across all splits, respectively.141414We verify that there is no leakage of dev and test splits of Fleurs in SeamlessM4Tv2-Large and Whisper-v3-Large. We observe that the averages and the standard deviations CER and sacreBLEU by task and language for either model are highly comparable across splits. We trust that the models have not been trained on the dev and test set of Fleurs. We ‘macro-average’ the metrics over all languages.
@@
S2ETT. SeamlessM4Tv2-Large outperforms Whisper-v3-large across the board in speech-to-English-text-translation and is favored for all but 2 languages. For both groups of languages supported by Whisper-v3-Large and SeamlessM4Tv2-Large, respectively, SeamlessM4Tv2-Large achieves on average about 8 higher sacreBLEU than Whisper-v3-Large. The gap from SeamlessM4T-Large to Whisper-v3-Large reduces to about 5 sacreBLEU for unsupported languages. This supports the notion that SeamlessM4Tv2-Large is a more robust multilingual speech encoder.
The more robust ASR and much stronger translation performance of SeamlessM4Tv2-Large results stem from its pre-training. SeamlessM4Tv2 first initializes a text encoder and decoder with weights from a pre-trained NLLB translation model (Team et al., 2022) and a speech encoder pre-trained on 4.5M hours of self-supervised training on w2v-BERT 2.0 objective. The model is then trained on translation objectives from text and speech to text between any two languages in both translation directions. The text encoder-decoder backbone is used to train the speech encoder with token-level knowledge distillation objectives on decoder output representations. On the contrary, Whisper-v3 trains models from scratch on, among others, in-language ASR and S2ETT (cf. §4). Consequently, the mixture of strong initialization from existing MT backbones, text-to-speech knowledge distillation, and multimodal translation objectives result in much stronger translation performance.
@@
Length Adaptor. The sequence length of encoded speech typically far exceeds the length of embedded tokenized text for the same input. The mSE of SeamlessM4Tv2-Large thus appends a temporal convolution, a ‘length adaptor’, as its final layer to reduce the resolution of speech frames by a factor of 8 and to better align the modalities (Seamless Communication et al., 2023).151515For architectural details, we refer to the original paper (Communication et al., 2023). This begs the question whether a length adaptor is essential for pooling semantics from speech output tokens. We hence compare SeamlessM4Tv2-Large’s performance on SIB-Fleurs with and without the length adaptor.
Table 4 presents the inconclusive results for both variants. For English, performance improves on high-quality utterances but decreases on lower-quality ones. In contrast, performance most pronouncedly declines for low-quality utterances in zs-xlt to supported languages (-2.6%). The zs-xlt unsupported languages is not affected for high-quality utterances (-0.1%) and only slightly for low-quality ones (-0.9%). Three factors may explain this modest gap. First, removing the length adaptor reduces model size by 47M parameters (-7.4 pp), which were part of substantial pre-training. Second, the length adaptor, as the final layer of the mSE, is most explicitly trained to embed speech into a shared multilingual space, attended to by the text decoder at every layer. Finally, the length adaptor might filter noisy frames through temporal downsampling to improve the robustness of speech encodings. While a modest gap persists across setups, we cannot decisively infer whether appending a length adaptor is, as opposed to replacing model capacity with other layers of equal parameter size during pre-training, crucial for improved multilingual SLU. We conclude that the pre-training regime is more vital for multilingual SLU capabilities of mSE than nuanced architectural design choices. We leave a more detailed investigation into the utility of length adaptors in SLU to future work.
ZS-XLT | |||||||||
EN (1) | S4T (89) | N/A (12) | AVG (101) | ||||||
S4Tv2-L | Size | B | W | B | W | B | W | B | W |
Incl. LA | 635M | 87.4 | 88.5 | 82.2 | 79.1 | 55.7 | 52.9 | 79.0 | 75.8 |
Excl. LA | 588M | 88.5 | 86.4 | 80.9 | 77.5 | 55.6 | 52.0 | 77.8 | 74.4 |
47M | +1.1 | -2.1 | -1.3 | -2.6 | -0.1 | -0.9 | -1.2 | -1.4 |